# Faraday Depolarization Study of a Radio Galaxy Using LOFAR Two-metre Sky Survey: Data Release 2

### v1

@@ 7 Summary and Conclusion
In this study, we have investigated the depolarization properties of a radio galaxy ILTJ012215.21+254334.8 using polarization data from the LOFAR Two-metre Sky Survey (LoTSS) DR2 catalogue. Polarization analysis provides direct insight into both the magnetic field structure and the magneto-ionic environment along the line of sight, offering a diagnostic tool for probing astrophysical jets, lobes, and the intervening medium.
@@
The necessity for multiple external depolarization components suggests a complex magneto-ionic medium with significant turbulence and inhomogeneities. Direct observational evidence for depolarization includes the contraction of the polarization vector radius in the - plane and a hint of monotonic decline in polarized fraction as a function of wavelength squared. Posterior analysis further supports the role of both turbulence and line-of-sight structure in shaping the observed polarization. Selecting a source with a relatively large projected linear size (335 kpc) has enabled a robust investigation of depolarization effects in a typical low-frequency radio galaxy. The polarized intensity, RM properties measured are consistent with the LoTSS DR2 sample.
In conclusion, our results demonstrate the efficacy of low-frequency polarization observations, combined with quantitative qu-fitting, in separating instrumental effects from astrophysical depolarization and Faraday rotation signatures. The best-fitting model reveals the importance of multi-component depolarization in a typical radio galaxies of the LoTSS DR2 catalogue, reflecting the inherent complexity of their magnetized environments, usually found in dense cluster environments. In future, extending this analysis to the full FR-I sample in the LoTSS DR2 catalogue, will help establish broader statistical trends and further constrain the interplay between turbulence, environmental density, source size, and polarization properties in extragalactic radio sources.
@@ Acknowledgments
SSP and AG would like to express their gratitude to Banwarilal Bhalotia College, Asansol, under Kazi Nazrul University, for facilitating the M.Sc. Physics program. This research forms part of the M.Sc. thesis work conducted in the final semester. The authors appreciate ChatGPT for helping with AI-driven copy editing and improving the manuscript’s clarity and style. AG acknowledges the financial support received through the SERB-SURE grant (SUR/2022/000595) from the Science and Engineering Research Board (SERB), a statutory body under the Department of Science and Technology (DST), Government of India. AG would also like to thank IUCAA, Pune, for their support through the Associateship Programme and for providing access to their computational facilities.
Appendix A Fit results for 1Tk+2ED model
In this appendix, we present the fit results for the next best-fit depolarization model, designated as the model. The following plots illustrate the polarization fraction (p, q, u), polarization angle as functions of wavelength squared, alongside the modelled curves derived from the best-fit parameters (Figure 5, and 6). We also include the posterior distribution of the model parameters which are displayed through one and two-dimensional projections in a corner plot (Figure 7).

### v2

@@ 7 Summary and Conclusion
In this study, we have investigated the depolarization properties of the radio galaxy ILTJ012215.21-254334.8 using polarization data from the LOFAR Two-metre Sky Survey (LoTSS) DR2 catalogue. An important outcome of this exercise is that, even using LoTSS HBA data alone, the Bayesian qu-fitting framework can discriminate between one-, two-, and three-component depolarization models and deliver well-defined posteriors for the dominant external dispersion components. This shows that LoTSS-DR2 survey data alone can robustly constrain multi-component Faraday depolarization models for typical sources in the DR2 catalog.
@@
The necessity for multiple external depolarization components suggests a complex magneto-ionic medium with significant turbulence and inhomogeneities. Direct observational evidence for depolarization is the monotonic decline in shown in Figure 2, which represents wavelength-dependent loss of coherence in the polarized signal. Posterior analysis further supports the role of both turbulence and line-of-sight structure in shaping the observed polarization. Selecting a source with a relatively large projected linear size (335 kpc) has enabled a robust investigation of depolarization effects in a typical low-frequency radio galaxy. The polarized intensity, RM properties measured are consistent with the LoTSS DR2 sample.
An important limitation of the present analysis is its intrinsic insensitivity to very Faraday-thick components. Because the LoTSS HBA band is confined to low frequencies, any emission with significantly larger than a few rad m-2 would be almost completely depolarized across the band and would not contribute to the observed spectra, even if intrinsically bright and highly polarized. The fitted external dispersion components therefore probe only the moderately Faraday-thin or thick part of the magneto-ionic environment; additional, much more heavily dispersed screens may be present but remain invisible to LoTSS and can only be constrained by combining these data with higher-frequency polarization measurements.
In conclusion, our results demonstrate the efficacy of low-frequency polarization observations, combined with quantitative qu-fitting, in separating instrumental effects from astrophysical depolarization and Faraday rotation signatures. The best-fitting model reveals the importance of multi-component depolarization in a typical radio galaxies of the LoTSS DR2 catalogue, reflecting the inherent complexity of their magnetized environments, usually found in dense cluster environments. In future, extending this analysis to the full FR-I sample in the LoTSS DR2 catalogue, will help establish broader statistical trends and further constrain the interplay between turbulence, environmental density, source size, and polarization properties in extragalactic radio sources. Thus, while this single-source study does not aim to provide a complete physical model of ILTJ012215.21-254334.8, it serves as a proof of concept for extracting robust Faraday-depolarization constraints from LoTSS-only data, and it quantifies both the power and the limitations of such an approach. The derived multi-component structure in Faraday depth, the required level of RM dispersion, and the model-selection behaviour provide concrete benchmarks for a forthcoming systematic analysis of the full FR-I subsample in the LoTSS-DR2 RM Grid.
@@ Acknowledgments
SSP would like to thank Banwarilal Bhalotia College, Asansol, under Kazi Nazrul University, for facilitating the M.Sc. Physics program. This research forms part of the M.Sc. thesis work conducted in the final semester. The authors appreciate ChatGPT for helping with AI-driven copy editing and improving the manuscript’s clarity and style. AG acknowledges the financial support received through the SERB-SURE grant (SUR/2022/000595) from the Science and Engineering Research Board (SERB), a statutory body under the Department of Science and Technology (DST), Government of India. AG would also like to thank IUCAA, Pune, for their support through the Associateship Programme and for providing access to their computational facilities. We would like to sincerely thank the reviewers for their thoughtful and constructive feedback on our manuscript.